OpenStudy (anonymous):

Write the expression as a complex number in standard form. Here is the problem 4i(3-i)(-2+8i)

OpenStudy (anonymous):

-104+8i

OpenStudy (saifoo.khan):

\[-104 + 8i\]

OpenStudy (anonymous):

I have to break the problem down I see how you got that answer. I just need to see it broken down.

OpenStudy (anonymous):

Okay. Use the distributive property. You have: \[4i(3-i)(-2+8i)\] Firstly distribute through the 4i: \[((3)(4i)-(i)(4i))(-2+8i)=(12i-4i^2)(-2+8i)\] But you know that: \[i^2=(\sqrt{-1})^2=-1\] So: \[(12i-4(-1))(-2+8i)=(12i+4)(-2+8i)\] Now, call: A=(12i+4) giving: \[A(-2+8i)\] Now distribute: \[-2A+8iA\] But A=12i+4: \[-2(12i+4)+8i(12i+4)=-24i-8+96i^2+32i=-24i-8-96+32i=-104+8i\]

OpenStudy (anonymous):

That a good explanation?

OpenStudy (anonymous):

Thank you malevolence19

OpenStudy (anonymous):

No problem xPPP